# The Dual-Route Model of Induction

### v1

@@ Abstract
Prior work on in-context copying has shown the existence of induction heads, which attend to and promote individual tokens during copying. In this work we introduce a new type of induction head: concept-level induction heads, which copy entire lexical units instead of individual tokens. Concept induction heads learn to attend to the ends of multi-token words throughout training, working in parallel with token-level induction heads to copy meaningful text. We show that these heads are responsible for semantic tasks like word-level translation, whereas token induction heads are vital for tasks that can only be done verbatim, like copying nonsense tokens. These two “routes” operate independently: in fact, we show that ablation of token induction heads causes models to paraphrase where they would otherwise copy verbatim. In light of these findings, we argue that although token induction heads are vital for specific tasks, concept induction heads may be more broadly relevant for in-context learning.
@@ 1 Introduction
LLMs can copy text using induction heads, a type of circuit found in decoder-only transformer models that enables them to copy sequences in-context (Elhage et al., 2021; Olsson et al., 2022). Given the sequence w|ax|wing...w, an induction head at the second occurrence of w would attend to and promote the earlier token ax by scanning for previous token information. Prior work has argued that induction heads are responsible for broader in-context learning capabilities (Olsson et al., 2022). However, it is unclear how attention heads operating on a token-by-token basis might handle “fuzzy” copying tasks like translation, which often requires conversion between words of differing token lengths (e.g., p|ommes| des| terre $\rightarrow$ pot|atoes). In this paper, we show that LLMs use concept induction heads in parallel with token induction heads to copy meaningful text.
Figure 1: The dual-route model of induction. LLMs develop token induction heads, which are used for verbatim copying, and concept induction heads, important for translation and “fuzzy” copying tasks. These two routes work in parallel to copy meaningful text.
@@
Concept Induction. Figure 1 also depicts our current understanding of concept induction heads, which are responsible for copying lexical information. In this work, we focus on concepts that take the form of lexical items (multi-token words and entities), as discussed by Feucht et al. (2024). To isolate these heads, we define a concept copying score based on causal mediation for a simple multi-token copying task (Section 2). Instead of attending to the next token, we find that these heads attend to the ends of multi-token words (Section 3), where concept information is more likely to be stored (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023; Kaplan et al., 2025). Concept induction heads are vital for semantic copying tasks (Section 4), and output word representations that are language-agnostic (Section 5).
@@ 2.1 Approach
This yields a clean prompt $p_{clean}=x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$, which we corrupt by re-sampling all of the tokens in the first half: $p_{corrupt}=y_{1}y_{2}...y_{n+\ell}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$. We patch the activations of each attention head $a^{(l,h)}$ (layer $l$, head $h$) from the penultimate token position of $p_{clean}$ (i.e., from $x^{\prime}_{n}$) into the same position in $p_{corrupt}$. Then, the concept copying score for head $(l,h)$ over concept set $\mathcal{C}$ is defined as
@@
For this experiment and for Section 3, we use the CounterFact dataset (Meng et al., 2022), which consists of subject-object relations (e.g., “Paul Chambers plays bass”). We sample a subset of subjects from this dataset to use as our set of concepts $\mathcal{C}$, where $|\mathcal{C}|=1024$. We could also use generic multi-token words, but results from Feucht et al. (2024) suggest that models treat both types of sequences similarly. We also measure scores for $P(c_{1})$ and $P(r_{2})$, but do not focus on them in this work (see Appendix A).
@@
Table 1: Models used in this paper. $|\mathcal{V}|$ is the model’s token vocabulary size, and $t$ is the number of tokens the model was trained on (in trillions).
@@ 3.1 Approach
To capture this attention behavior, we design a last-token matching score which measures how much attention is paid to the last token of a multi-token concept. First, we select a concept $c=c_{1}...c_{m}$ from CounterFact subjects $\mathcal{C}$, evenly sampling across concept lengths $2\leq m\leq 5$. We then construct random repeated sequences of tokens as before (see §2) and insert the concept on the left to create the prompt $x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}$. The last-token matching score is then calculated as an average over the concept set $\mathcal{C}$:
@@ 4.2 Results
Box 4.2: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [bar[i] for i in range( len(bar)) if i % 2 == 0]
@@ 5 Concept Induction is Language-Agnostic
If concept induction heads are important for translation tasks, what do they output? We hypothesize that the representations being copied by concept induction heads are abstract representations of word meanings. In other words, we posit that concept induction heads have the same activations when copying “waxwing” as they do when copying “свиристель,” as these two words refer to the same concept, and are only expressed differently on a surface level. This is in line with previous work suggesting that concept information in LLMs may be language-agnostic (Wendler et al., 2024; Dumas et al., 2024; Brinkmann et al., 2025), though Schut et al. (2025) argues concepts for some tasks may be biased towards English.
@@ 5.2 Results
Figure 6 shows results for Llama-2-7b when patching outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. Patching concept induction heads causes the model to output the source word $w_{s}$ in the base language $\smash{\ell^{(out)}_{b}}$. For example, patching the outputs of concept heads from a Spanish-Italian prompt translating “child” into a Japanese-Chinese prompt causes the model to output the word for “child” in Chinese. This intervention is the most effective at about $k=80$, which also corresponds to the point where translation and nonsense copying are the most separate in Figure 5c. We show similar results for Llama-3-8b and more language pairs in Appendix E. These experiments suggest that concept induction heads transport semantic representations that are expressible in multiple languages.
Figure 6: (a) Patching the top-$k$ concept induction head outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. (b) Patching concept induction heads changes the concept output by the model, without affecting the language (i.e. the Spanish word is output in Chinese instead of Italian). This effect is strongest for $k=80$, which is also where the largest separation between semantic and literal copying performance is in Figure 5c.
6 Function Vector Heads Versus Concept Induction Heads
Figure 7: (a) Patching concept heads changes the output meaning without affecting language (Figure 6), while patching FV heads changes the output language without affecting meaning. (b) Patching FV heads from a French-English prompt into a German-Russian prompt ($n=128$) flips the output language to English (green) without changing the concept (red).
Function vector (FV) heads are attention heads whose outputs help to promote in-context tasks; e.g., the outputs of FV heads for a few-shot antonym task can also cause the model to output antonyms in new contexts (Todd et al., 2024). We notice that the highest-scoring concept copier head is also the highest-scoring FV head by a large margin for both Llama models. We calculate correlations between FV scores and concept copying scores and find significant, albeit weak, positive correlations for all models, suggesting there may be some relationship between these two types of heads (Figure 38).
@@ 6 Function Vector Heads Versus Concept Induction Heads
However, while FV and concept induction heads seem to be related, they seem to play two distinct roles in semantic copying tasks. We focus on translation, and patch outputs of FV heads between translation prompts using the same approach that we do for concept induction heads in Section 5. Our experimental setup and data is identical, except we patch activations $\smash{a_{s}^{k}=\{a_{s}^{(l,h)}|(l,h)\in F^{k}\}}$, where $\smash{F^{k}}$ represents the top-$k$ FV heads for a given model. We then measure the output of the base word in the source output language, $\smash{w_{b}:\ell_{s}^{(out)}}$. Figure 7 shows these results for Llama-2-7b. For the exact same example, FV heads cause the model to output the same concept in a different language, whereas concept induction heads cause the model to output a different concept in the same language.
@@
We also replicate ablations from Section 4 using FV rankings instead of concept and token copying rankings (Appendix F). Ablation of FV heads damages performance for non-verbatim tasks significantly, which makes sense: models cannot perform a task without knowing what the task is.
@@ 8 Conclusion
Acknowledgments
Thanks to Chris Wendler, Clément Dumas, and Kerem Sahin for helpful discussions on multilingual representations. We thank Hiba Ahsan and Can Rager for comments on our draft. SF also thanks Adrian Chang, Jerry H. Wang, and Emily Wang for feedback throughout the process of writing this paper. We thank the Center for AI Safety for providing computational resources used for certain experiments. Experiments are implemented in NNsight (Fiotto-Kaufman et al., 2025). SF, ET, and DB are funded by a grant from Open Philanthropy.
Ethics Statement
In this work, we study models trained primarily on English text. Although we claim that these models have representations that are agnostic to language, we acknowledge that these representations are likely biased towards English. In other words, they may disproportionately reflect the perspectives of people from English-speaking regions.
@@ Appendix B Token and Concept Induction Heads Throughout Training
We include analysis of causal and attention-based induction scores over time for OLMo-2-7b and Pythia-6.9b (models that provide access to training checkpoints). In Figures 17 and 19, we show token and concept copying scores from Section 2 over checkpoints. Figures 18 and 20 show next-token and last-token attention scores from Section 3 over checkpoints.
@@ C.1 Attention Scores for Top 64 Heads
We include expanded versions of Figure 4 for the top-64 token and concept copier heads for each model. See Figures 21 through 28. We see that concept copier heads tend to have high last-token matching scores (§3), but also that each model has a few heads that seem to do both token and concept induction. Heads that appear in the top-64 of both token-copying scores and concept-copying scores are bolded.
@@ D.1 Full Results
We show the results of ablating token copier heads and concept copier heads on our “vocabulary list" tasks for all four models in Figures 31-34. We also show the number of heads that overlap for increasing choices of $k$ in Figure 30. Mean-ablation is done using average activations of all heads over $n=1024$ Pile (Gao et al., 2020) documents.
@@
Figure 34: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Pythia-6.9b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. Unlike previous models, semantic tasks (translation, synonyms, and antonyms) are also damaged when we ablate token induction heads (left). However, capitalization and English copying accuracy remains relatively high. We do not have a good explanation for this result. Because Pythia is trained on an order of magnitude fewer tokens than other models in this work (0.3 trillion, versus $\geq$ 2 trillion), it may be that concept induction and token induction are somewhat fused.
@@ D.2 Examples of Prompts
We show examples of two semantic task prompts from Section 4 in Box D.2. We also show examples of two verbatim tasks in Box D.2.
Box D.1: Examples of Vocabulary List Prompts (French-English & Antonym) <s> French vocab:  1. completement  2. inconstitutionnalité  3. flocon  4. chiens  5. vaccinés  6. racontée  7. spécialités  8. parfumée  9. condensateur  10. saumons  English translation:  1. completely  2. unconstitutionality  3. flake  4. dogs  5. inoculated  6. told  7. specialties  8. fragrant  9. capacitor  10. <s> Words:  1. hydrous  2. enlarge  3. alternative  4. messy  5. traditionalism  6. simplicity  7. underwater  8. transitive  9. omit  10. contract  Antonyms:  1. anhydrous  2. cut  3. primary  4. neat  5. rationalism  6. difficulty  7. amphibious  8. intransitive  9. remember  10.
Box D.2: Examples of Vocabulary List Prompts (English & Nonsense Copying) <s> English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. recover  English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. <s> Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10. ercase  Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10.

### v2

@@ Abstract
Prior work on in-context copying has shown the existence of induction heads, which attend to and promote individual tokens during copying. In this work we discover a new type of induction head: concept-level induction heads, which copy entire lexical units instead of individual tokens. Concept induction heads learn to attend to the ends of multi-token words throughout training, working in parallel with token-level induction heads to copy meaningful text. We show that these heads are responsible for semantic tasks like word-level translation, whereas token induction heads are vital for tasks that can only be done verbatim (like copying nonsense tokens). These two “routes” operate independently: we show that ablation of token induction heads causes models to paraphrase where they would otherwise copy verbatim. By patching concept induction head outputs, we find that they contain language-independent word representations that mediate natural language translation, suggesting that LLMs represent abstract word meanings independent of language or form.
@@ 1 Introduction
LLMs can copy text using induction heads, a type of circuit found in decoder-only transformer models that enables them to copy sequences in-context (Elhage et al., 2021; Olsson et al., 2022). Given the sequence w|ax|wing...w, an induction head at the second occurrence of w would attend to and promote the earlier token ax by scanning for previous token information. Prior work has argued that induction heads are responsible for broader in-context learning capabilities (Olsson et al., 2022). However, it is unclear how attention heads operating on a token-by-token basis might handle “fuzzy” copying tasks like translation, which often requires conversion between words of differing token lengths (e.g., p|ommes| de| terre $\rightarrow$ pot|atoes). In this paper, we show that LLMs use concept induction heads in parallel with token induction heads to copy meaningful text.
Figure 1: The dual-route model of induction. LLMs develop token induction heads, which are used for verbatim copying, alongside concept induction heads, important for translation and “fuzzy” copying tasks. These two routes work in parallel to copy meaningful text.
@@
Concept Induction. Figure 1 also depicts our current understanding of concept induction heads, which are responsible for copying lexical information. To isolate these heads, we define a concept copying score based on causal mediation for a simple multi-token copying task (Section 2). Instead of attending to the next token, we find that these heads attend to the ends of multi-token words (Section 3), where concept information is more likely to be stored (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023; Feucht et al., 2024; Kaplan et al., 2025). Concept induction heads are vital for semantic copying tasks (Section 4), and output word representations that are language-agnostic (Section 6).
@@ 2.1 Approach
This yields a clean prompt $p_{clean}=x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$, which we corrupt by replacing the first half with different random tokens: $p_{corrupt}=y_{1}y_{2}...y_{n+m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$. We patch the activations of each attention head $a^{(l,h)}$ (layer $l$, head $h$) from the penultimate token position of $p_{clean}$ (i.e., from $x^{\prime}_{n}$) into the same position in $p_{corrupt}$. Then, the concept copying score for head $(l,h)$ over concept set $\mathcal{C}$ is defined as
@@
For this experiment and for Section 3, we take concepts from the CounterFact dataset (Meng et al., 2022), which consists of subject-object relations (e.g., “Paul Chambers plays bass”). We sample a subset of subjects from this dataset to use as our set of concepts $\mathcal{C}$, where $|\mathcal{C}|=1024$. We could also use generic multi-token words, but results from Feucht et al. (2024) suggest that models treat both types of sequences similarly. We also measure scores for $P(c_{1})$ and $P(r_{2})$, but do not focus on them in this work (see Appendix A).
@@
Table 1: Models used in this paper. $|\mathcal{V}|$ is the model’s token vocabulary size, and $t$ is the number of tokens the model was trained on (in trillions). We evaluate OLMo-2-1b on only a subset of experiments.
@@ 3.1 Approach
To capture this attention behavior, we design a last-token matching score which measures how much attention is paid to the last token of a multi-token concept. First, we select a concept $c=c_{1}...c_{m}$ from CounterFact subjects $\mathcal{C}$, evenly sampling across concept lengths $2\leq m\leq 5$. We then construct random repeated sequences of tokens as before (see §2) and insert the concept at the end of the first half to create the prompt $x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}$. The last-token matching score is then calculated as an average over the concept set $\mathcal{C}$:
@@ 4.2 Results
Box 4.2: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         asdf.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [bar[i] for i in range( len(bar)) if i % 2 == 0]
5 Concept Heads Reveal Semantics of Hidden States
If concept induction heads work with meaningful representations, the subspaces that they read from and write to must contain semantic information. What do these subspaces look like? We use the attention weights of concept heads to define a concept lens $\smash{L_{C_{k}}\in\mathbb{R}^{(d,d)}}$ that reveals the semantic information contained within any particular hidden state. Specifically, we sum the OV matrices (Elhage et al., 2021) of the top-$k$ concept induction heads $C_{k}$ to obtain $L_{C_{k}}$. We then apply this transformation to a hidden state, followed by the model’s final normalization and decoding head (see Appendix E for details).
We can use this linear transformation to visualize how a model represents a word in a particular context. For example, consider the word cardinals, which could refer to a sports team, senior members of the Catholic church, or a group of red birds. By transforming the hidden state for the token inals with $L_{C_{k}}$ and projecting to vocabulary space, we can see that Llama-2-7b has a different semantic representation for cardinals depending on the context that the word is in, further supporting the idea that concept heads represent word meanings rather than surface-level token information.
Prompt Concept Lens Outputs ($l=20$)
cardinals ’ Card’, ’ cardinal’, ’Card’
he was a lifelong fan of the cardinals ’ Card’, ’ football’, ’ baseball’,
the secret meeting of the cardinals ’ Rome’, ’ Catholic’, ’ Card’
in the morning air, she heard northern cardinals ’ birds’, ’ bird’, ’ Bird’,
Table 2: Applying concept lens to hidden states for the token inals in the word cardinals reveals context-sensitive semantic representations. We show layer $l=20$ with weights from the top $k=80$ concept induction heads for Llama-2-7b. See Appendix E.
@@ 6 Concept Induction is Language-Agnostic
Figure 6: (a) Patching the top-$k$ concept induction head outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. (b) Patching concept induction heads changes the concept output by the model without affecting the language (i.e., niño, the Spanish word for “child,” is translated into Chinese instead of Italian). This effect is strongest for $k=80$, which is also where the largest separation between semantic and literal copying performance is found in Figure 5c. Across $n=128$ examples, this approach causes the model to output the source Spanish word in the base output language with an accuracy of about 0.40 (solid red line). This is comparable to the model’s original Japanese-Chinese translation accuracy of 0.48 (dotted gray line).
If concept induction heads are important for translation tasks, what do they output? We hypothesize that the representations being copied by concept induction heads are abstract representations of word meanings. In other words, we posit that concept induction heads have the same activations when copying “waxwing” as they do when copying “свиристель,” as these two words refer to the same concept, and are only expressed differently on a surface level. This is in line with previous work suggesting that concept information in LLMs may be language-agnostic (Wendler et al., 2024; Dumas et al., 2024; Brinkmann et al., 2025), though Schut et al. (2025) argues concepts for some tasks may be biased towards English. For model representations to be truly language-agnostic, they must be trained on those languages—thus, even if high-resource languages are represented in a unified semantic space, this effect may not hold for low-resource languages. Unfortunately, as we do not evaluate on low-resource languages in this work, we cannot make claims as to how models represent low-resource languages.
@@ 6.2 Results
Figure 6 shows results for Llama-2-7b when patching outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. Patching concept induction heads causes the model to output the source word $w_{s}$ in the base language $\smash{\ell^{(out)}_{b}}$. For example, patching the outputs of concept heads from a Spanish-Italian prompt translating “child” into a Japanese-Chinese prompt causes the model to output the word for “child” in Chinese. This intervention is the most effective at about $k=80$, which also corresponds to the point where translation and nonsense copying are the most separate in Figure 5c. We show similar results for Llama-3-8b and more language pairs in Appendix F. This suggests that concept induction heads transport semantic representations that are expressible across multiple languages.
7 Function Vector Heads Complement Concept Induction Heads
Figure 7: (a) Patching concept heads changes semantics without affecting language, while patching FV heads changes output language without affecting meaning. In red, we show the same experiment from Figure 6: here, patching concept heads causes the model to output the source concept “committee” in Russian. In green, we show that patching FV heads at the same position for the same prompt causes the model to output the base concept “toilet” in English. (b) Across $n=128$ examples, patching FV heads from a French-English prompt into a German-Russian prompt flips the output language to English with an accuracy of about 0.80 (green line). This is higher than the model’s original German-Russian translation accuracy (gray dotted line, approximately 0.41), perhaps because translating into English is easier for Llama-2-7b than translating into Russian.
Function vector (FV) heads are attention heads whose outputs help to promote in-context tasks. The outputs of FV heads for a few-shot antonym task, for example, will cause the model to output antonyms when added to new contexts (Todd et al., 2024). We argue that FV heads can be thought of as “sisters” to concept induction heads: in the case of translation, concept heads copy semantic information, whereas FV heads copy language information.
We calculate correlations between FV scores and concept copying scores and find significant, albeit weak, positive correlations for all models (Figure 48), suggesting there may be some relationship between these two types of heads. This is perhaps because they can both be thought of as “soft” induction heads, responsible for copying high-level conceptual information in-context.
@@ 7 Function Vector Heads Complement Concept Induction Heads
However, FV and concept heads seem to play two distinct roles in the tasks we examine. We focus on translation, and patch outputs of FV heads between translation prompts using the same approach that we do for concept induction heads in Section 6. Our experimental setup and data is identical, except we patch activations $\smash{a_{s}^{k}=\{a_{s}^{(l,h)}|(l,h)\in F^{k}\}}$, where $\smash{F^{k}}$ represents the top-$k$ FV heads for a given model. We then measure the output of the base word in the source output language, $\smash{w_{b}:\ell_{s}^{(out)}}$. In other words, we measure whether patching FV heads changes the output language while retaining the original base concept.
Figure 7 shows that for the exact same prompts, patching FV heads causes the model to output the same concept in a different language, whereas patching concept induction heads causes the model to output a different concept in the same language. The outcome of patching FV heads is more sensitive to output language (i.e., the effect is strongest when the output language is English), but nonetheless suggests that FV heads play a distinct role from concept induction heads (see Appendix G).
@@
We also replicate ablations from Section 4 using FV rankings instead of concept and token copying rankings (Appendix G). Ablation of FV heads damages performance for non-verbatim tasks significantly, which makes sense: models cannot perform a task without knowing what the task is.
@@ Appendix A Token & Concept Copying Scores
Figure 15: Probability differences when patching head activations for OLMo-2-1b over entity tokens. The right-hand side shows concept copying scores; we do not utilize the left-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $c_{1}$ at that token position (left), and increase in probability for $c_{2}$ at the following token position. See Figure 2 and Section 2 for details. Unlike other models in this paper, OLMo-2-1b has only 16 layers and 16 heads per layer.
Figure 16: Probability differences when patching head activations for OLMo-2-1b over random tokens. The left-hand side shows token copying scores; we do not utilize the right-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $r_{1}$ at that token position (left), and increase in probability for $r_{2}$ at the following token position. See Figure 2 and Section 2 for details. Unlike other models in this paper, OLMo-2-1b has only 16 layers and 16 heads per layer.
@@ Appendix B Token and Concept Induction Heads Throughout Training
We include analysis of causal and attention-based induction scores over time for OLMo-2-7b and Pythia-6.9b (models that provide access to training checkpoints). In Figures 19 and 21, we show token and concept copying scores from Section 2 over checkpoints. Figures 20 and 22 show next-token and last-token attention scores from Section 3 over checkpoints.
@@ C.1 Attention Scores for Top 64 Heads
We include expanded versions of Figure 4 for the top-64 token and concept copier heads for each model. See Figures 23 through 30. We see that concept copier heads tend to have high last-token matching scores (§3), but also that each model has a few heads that seem to do both token and concept induction. Heads that appear in the top-64 of both token-copying scores and concept-copying scores are bolded.
@@ D.1 Full Results
We show the results of ablating token copier heads and concept copier heads on our “vocabulary list" tasks for all four models in Figures 33-37. We also show the number of heads that overlap for increasing choices of $k$ in Figure 32.
@@
Figure 36: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in OLMo-2-1b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. To be consistent with larger models, we show approximately the top 30% of heads.
Figure 37: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Pythia-6.9b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. Unlike previous models, semantic tasks (translation, synonyms, and antonyms) are also damaged when we ablate token induction heads (left). However, capitalization and English copying accuracy remains relatively high. Because Pythia is trained on an order of magnitude fewer tokens than other models in this work (0.3 trillion, versus $\geq$ 2 trillion), it may be that concept induction and token induction are somewhat fused.
@@ D.2 Examples of Prompts
We show examples of two semantic task prompts from Section 4 in red boxes below. We also show examples of verbatim tasks in blue.
Vocabulary List Prompt Example (French-English) <s> French vocab:   1. completement   2. inconstitutionnalité   3. flocon   4. chiens   5. vaccinés   6. racontée   7. spécialités   8. parfumée   9. condensateur   10. saumons English translation:   1. completely   2. unconstitutionality   3. flake   4. dogs   5. inoculated   6. told   7. specialties   8. fragrant   9. capacitor   10. Vocabulary List Prompt Example (Antonym) <s> Words:   1. hydrous   2. enlarge   3. alternative   4. messy   5. traditionalism   6. simplicity   7. underwater   8. transitive   9. omit   10. contract Antonyms:   1. anhydrous   2. cut   3. primary   4. neat   5. rationalism   6. difficulty   7. amphibious   8. intransitive   9. remember   10.
Vocabulary List Prompt Example (English) <s> English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. recover  English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. Vocabulary List Prompt Example (Nonsense Copying) <s> Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10. ercase  Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10.